# Hyper-Transforming Latent Diffusion Models

### v2

@@ Abstract
Keywords Diffusion Models, Implicit Neural Representations, Generative Modeling, Variational Inference
@@ 1 Introduction
Generative modelling has seen remarkable advances in recent years, with diffusion models achieving state-of-the-art performance across multiple domains, including images, videos, and 3D synthesis Ho et al. (2020); Dhariwal and Nichol (2021); Rombach et al. (2022). A key limitation of existing generative frameworks, however, is their reliance on structured output representations, such as pixel grids, which constrain resolution and generalisation across data modalities. In contrast, Implicit Neural Representations (INRs) have emerged as a powerful alternative that parametrises signals as continuous functions Sitzmann et al. (2020); Mildenhall et al. (2021). By leveraging INRs, generative models can represent complex data distributions at arbitrary resolutions, yet learning expressive distributions over INR parameters remains a fundamental challenge.
A common approach to modelling distributions over INRs is to use hypernetworks Ha et al. (2017), which generate the weights and biases of an INR conditioned on a latent code Dupont et al. (2022a); Koyuncu et al. (2023). However, MLP-based hypernetworks suffer from scalability bottlenecks, particularly when generating high-dimensional INRs, as direct parameter regression constrains flexibility and expressiveness. More recently, Transformer-based hypernetworks have been proposed to alleviate these issues, introducing attention mechanisms to efficiently predict INR parameters Chen and Wang (2022); Zhmoginov et al. (2022). Nonetheless, existing approaches such as Trans-INR Chen and Wang (2022) remain deterministic, limiting their applicability in probabilistic frameworks.
In this work, we introduce a novel framework for INR generation, named Latent Diffusion Models of INRs (LDMI), that incorporates our proposed Hyper-Transformer Decoder (HD), a probabilistic Transformer-based decoder for learning distributions over INR parameters. Our approach combines the strengths of hypernetworks with recent advances in meta-learning for INRs while integrating into latent diffusion-based generative frameworks. Unlike previous works, the Hyper-Transformer employs a full Transformer architecture, where a Transformer Encoder processes latent variables, and a Transformer Decoder generates INR parameters via cross-attention. This design enables flexible, probabilistic generation of neural representations across diverse data modalities, surpassing the deterministic limitations of prior Transformer-based hypernetworks Chen and Wang (2022); Zhmoginov et al. (2022).
The Hyper-Transformer supports two training paradigms: (i) full training, where the model is trained from scratch alongside a latent diffusion model (LDM; Rombach et al., 2022), and (ii) hyper-transforming, where a pre-trained LDM is adapted by replacing its decoder with the Hyper-Transformer, allowing for efficient transfer learning without retraining the entire generative pipeline. This flexibility enables the Hyper-Transformer to scale effectively while leveraging existing pre-trained diffusion models.
@@
Figure 1: Uncurated samples from LDMI at multiple resolutions.
@@ 2.1 Latent Diffusion Models
First stage The first stage involves learning the parameters of an encoder $\mathcal{E}_{\psi}(\cdot)$ and a decoder $\mathcal{D}_{\lambda}(\cdot)$, such that accurate likelihoods $p_{\lambda}({\bm{y}}|{\bm{z}})$ are achieved via meaningful, highly structured latent representations $q_{\psi}({\bm{z}}|{\bm{y}})$. Both the $\beta$-VAE Higgins et al. (2017) and Vector-Quantized VAE (VQ-VAE; Van Den Oord et al., 2017) are valid choices for learning the structured latent space in the first stage. In this stage, especially for image data, perceptual losses can be incorporated to further improve the quality of the generated images Hou et al. (2017); Dosovitskiy and Brox (2016); Hou et al. (2019). In this study, we leverage both the $\beta$-VAE and VQ-VAE approaches, selecting the most suitable configuration depending on the dataset characteristics.
Second stage To bridge the gap between a highly structured aggregate posterior and an overly simplistic standard prior, and to achieve high-quality unconditional generations, the second stage learns an expressive prior that approximates the learned aggregate posterior $q_{\psi}({\bm{z}})$ by fitting a diffusion model $p_{\theta}({\bm{z}})$. While both the continuous-time SDE-based approach Song et al. (2021a) and the discrete Markov chain-based formulation Ho et al. (2020) are viable, in this work, we follow the DDPM-based approach since the original LDM implementation Rombach et al. (2022) adopts the Denoising Diffusion Probabilistic Model (DDPM; Ho et al., 2020) framework for training. Within this framework, like the most successful models Saharia et al. (2022); Dhariwal and Nichol (2021), a reweighted variant of the variational lower bound on $q({\bm{z}})=\mathbb{E}_{p({\bm{y}})}\left[q({\bm{z}}|{\bm{y}})\right]$ is considered, which mirrors denoising score-matching Song et al. (2021a):
$L_{\text{DDPM}}=\mathbb{E}_{{\bm{y}},{\bm{z}},t,\epsilon}\left[\|\epsilon-\epsilon_{\theta}({\bm{z}}_{t},t)\|^{2}\right]$ (1)
where ${\bm{y}}\sim p({\bm{y}})$, ${\bm{z}}\sim q_{\psi}({\bm{z}}|{\bm{y}})$, $t\sim U(1,T)$ and $\epsilon\sim\mathcal{N}(\bm{0},\bm{I})$.
@@
In DDPM, the reverse posterior density is no longer Markovian and coincides with the inference model proposed later in DDIM Song et al. (2021b). In DDIM, it is demonstrated that faster sampling can be achieved without retraining, simply by using the posterior approximation for the estimation $\hat{{\bm{z}}}$, which re-defines the generative process as:
@@ 2.2 Probabilistic Implicit Neural Representations
Recent advances in Implicit Neural Representations (INRs; Sitzmann et al., 2020) have demonstrated their potential as a powerful method for directly parametrising continuous functions by mapping coordinates $({\bm{x}})$ to output features $({\bm{y}})$ using a neural network $f_{\Phi}({\bm{y}}|{\bm{x}})$ Mildenhall et al. (2021); Mescheder et al. (2019); Tancik et al. (2020). More recently, deep generative frameworks have been adapted to handling INRs, such that flexible conditional generation at arbitrary resolutions can be achieved Dupont et al. (2022b); Koyuncu et al. (2023). Unlike traditional generative models that aim to maximize the likelihood of structured data (e.g., fitting a VAE to approximate $p({\bm{y}})$), INRs offer a more flexible approach by approximating the conditional distribution $p({\bm{y}}|\bm{x})$. This flexibility allows for sampling data or measuring uncertainty at any desired location, providing advantages in cases where spatial or feature-based variability is essential.
With the aim to generate INRs, typically, a model meta-learns hidden representations for later transforming them into the parameters $\Phi$ of the INR. However, modelling these representations probabilistically is a key challenge, with two critical objectives: (i) embedding observed data into a posterior distribution to allow for conditional generation, and (ii) generating synthetic data by sampling from the prior distribution of hidden representations. Koyuncu et al. (2023) addressed these objectives by defining a latent variable ${\bm{z}}$ and proposing a variational framework that introduces a flow-based learnable prior $p_{\theta}({\bm{z}})$ and learns an approximate posterior via an encoder $q_{\psi}({\bm{z}}|{\bm{x}},{\bm{y}})$. To map ${\bm{z}}$ to the INR parameters, they proposed an MLP-based hypernetwork Ha et al. (2017). Consequently, scaling this approach to large, high-quality images proved difficult due to the complexities of the high-dimensional weight and bias space required for INRs, which necessitate a highly flexible prior and pose significant optimization challenges. Similarly, Dupont et al. (2022a) proposed adversarial methods, where the generator makes use of similar hypernetworks. While effective in mitigating adversarial training issues, this approach lacks support for conditional generation, a crucial component for many downstream tasks, and shares the bottleneck issues of the MLP-based hypernetwork.
More recently, two approaches have been proposed to mitigate the aforementioned issues. The first approach, introduced in Functa (Dupont et al., 2022b) and consecutive work (Bauer et al., 2023), divides the process into two stages: first, gradient-based meta-learn the neural fields per image by optimization, and second, training a deep generative model on the so-called functaset. The second approach, proposed by Chen et al. (2024) draws inspiration from LDMs Rombach et al. (2022) to pre-train a minimally-regularized autoencoder via $\beta$-VAE Higgins et al. (2017) or a VQ-VAE Van Den Oord et al. (2017) for later training a diffusion model on the learned posterior distribution.
@@ 2.3 Hyper-Transformers
A widely used approach within this framework is to employ hypernetworks (Ha et al., 2017) to map hidden representations $\bm{z}$ of a datapoint into parameters $\Phi=g_{\phi}({\bm{z}})$, where $g_{\phi}$ represents the hypernetwork with learnable parameters $\phi$, and $\Phi$ are the parameters modulating the INR generator ${\bm{y}}=f_{\Phi}({\bm{y}}|\bm{x})$. The hypernetwork shares parameters across the dataset, while the INR has parameters unique to each datapoint, generated dynamically by the hypernetwork. However, a significant challenge arises as the size of the INR network increases: directly generating all INR parameters as a flattened vector introduces bottlenecks in the hypernetwork’s capacity.
To overcome the limitations of MLP-based hypernetworks, recent advances have introduced Hyper-Transformers (Chen and Wang, 2022; Zhmoginov et al., 2022), which extend the capabilities of traditional hypernetworks by leveraging the attention mechanisms and scalability of Transformer architectures and generate INR parameters in modular chunks. This chunked approach allows the model to focus on relevant parameter subsets when generating weights and biases, avoiding the computational bottleneck of processing large flattened parameter vectors while maintaining network scalability. Building on this paradigm, recent work (Ruiz et al., 2024) introduces HyperDreamBooth, which similarly employs transformer-based hypernetworks but for generating low-rank adaptation parameters (LoRA) to personalize text-to-image diffusion models from single examples. While HyperDreamBooth focuses on efficient adaptation of existing generative models for personalization tasks, our LDMI framework extends the hyper-transformer approach to develop a complete generative model over the entire space of implicit neural representations. Our Hyper-Transformer Decoder generates full INR parameter sets enabling resolution-agnostic generation across diverse modalities (images, 3D fields, climate data), representing a distinct contribution to the emerging field of function-level generative modeling.
However, while these works have demonstrated impressive results in deterministic image reconstruction tasks, they do not yet extend to probabilistic frameworks that enable generation of synthetic data. To our knowledge, none of these architectures have been proposed as decoders of a latent variable model. In this work, we open a new promising direction by proposing full Transformers that are fed with tensor-shaped latent variables.
@@ 2.4 Function-valued stochastic processes
Several recent approaches have explored alternative methods for modeling function-valued stochastic processes. Neural Diffusion Processes (NDPs; Dutordoir et al., 2023) leverage a diffusion model to define distributions over function values at given coordinates. Their architecture explicitly enforces exchangeability and permutation invariance via a bi-dimensional attention mechanism, and their sampling mechanism mimics Gaussian processes and related meta-learning methods such as Neural Processes. Importantly, unlike our approach, the function itself is not represented via a neural network whose parameters are generated or learned—rather, the model learns to denoise function values directly, conditioned on inputs. Similarly, Simformer (Gloeckler et al., 2024) is designed for simulation-based inference (SBI), where the goal is to infer unknown parameters of stochastic simulators from observations. It treats both data and parameters as random variables and learns a diffusion model over the joint distribution $p(\bm{x},\bm{\theta})$, allowing for flexible sampling of any conditional (e.g., posterior, likelihood, marginals). While parameters may include function-valued (infinite-dimensional) components, they are not represented as INRs—rather, they are input variables within the inference pipeline. In contrast, our approach focuses on generating complete INR parameters via hyper-transformers, enabling a more direct representation of continuous functions across various modalities while maintaining resolution independence.
@@ 3 Methodology
(a) Inference model (top) and diffusion prior (bottom).
@@
Figure 2: Overview of the proposed method, LDMI. (a) The latent space models: the top illustrates the inference model, trained in the first stage, where an encoder maps data into the variational parameters of the approximate posterior; the bottom shows the DDIM prior, learned in the second stage, to approximate this posterior. (b) The full decoder: the HD module transforms the latent variable into the weights and biases of an INR, enabling continuous signal representation.
We introduce Latent Diffusion Models of Implicit Neural Representations (LDMI), a novel family of generative models operating in function space. While previous probabilistic INR approaches have explored latent variable models for INRs Koyuncu et al. (2023); Dupont et al. (2022a) and two-stage training schemes incorporating diffusion-based priors in the second stage Park et al. (2024); Dupont et al. (2022b); Bauer et al. (2023), these methods face key limitations. Notably, MLP-based hypernetworks suffer from capacity bottlenecks, while existing frameworks lack a unified and compact approach for handling diverse data modalities.
To address these challenges, we propose the Hyper-Transformer Decoder (HD), which extends the flexibility of hypernetworks with the scalability of Transformer-based architectures. Unlike prior deterministic frameworks Chen and Wang (2022), HD introduces a probabilistic formulation, enabling uncertainty modelling and improved generative capacity for INRs.
@@ 3.2 Generative Model
A common approach in prior works Koyuncu et al. (2023); Dupont et al. (2022a, b) is to model $\Phi$ implicitly using an auxiliary neural network known as a hypernetwork Ha et al. (2017), typically MLP-based, which modulates the INR parameters based on a lower-dimensional latent representation:
@@
where ${\bm{z}}\in\mathbb{R}^{H_{z}\times W_{z}\times d_{z}}$ is a tensor-shaped latent code with spatial dimensions $H_{z}\times W_{z}$ and channel dimension $d_{z}$. This latent space serves as a compressed representation of the continuous signal. However, MLP-based hypernetworks introduce a fundamental bottleneck: the final layer must output all parameters of the target INR, which leads to scalability issues when modulating high-capacity INRs. As a result, prior works employing MLP-based hypernetworks typically use small INR architectures, often restricted to three-layer MLPs Koyuncu et al. (2023); Dupont et al. (2022a). Additionally, these works rely on standard MLPs with ReLU or similar activations and often preprocess coordinates using Random Fourier Features (RFF) Tancik et al. (2020) to capture high-frequency details. However, RFF-based embeddings struggle to generalize to unseen coordinates without careful validation. In contrast, SIREN inherently captures the full frequency spectrum, making it more effective for super-resolution tasks.
However, using SIREN as the INR module introduces further challenges. As commented by their authors Sitzmann et al. (2020), optimizing SIRENs with not carefully chosen uniformly distributed weights yields poor performance both in accuracy and in convergence speed. This issue worsens when the weights are not optimized, but generated by a hypernetwork.
To overcome these limitations, we introduce the Hyper-Transformer Decoder (HD), a Transformer-based hypernetwork designed to scale effectively while preserving the flexibility of larger INRs and ensuring proper modulation of SIREN weights. The details of our approach are described in the following section.
Figure 3: Diagram of the Hyper-Transformer Decoder (HD). The latent variable ${\bm{z}}$ is tokenized and processed by a Transformer Encoder. A Transformer Decoder, initialized with learnable grouped weights $\bar{{\mathbf{W}}}_{l}^{\text{i}}$, cross-attends to the latent tokens to generate the set of grouped weights $\bar{{\mathbf{W}}}_{l}^{\text{o}}$. The full weight matrices ${\mathbf{W}}_{l}^{\text{o}}$ are then reconstructed by combining the grouped weights with learnable template weights $\bar{{\mathbf{W}}}_{l}^{\text{b}}$. Biases $\mathbf{b}_{l}$ are learned as global parameters.

### v3

@@ Abstract
Diffusion Models, Implicit Neural Representations, Generative Modeling, Variational Inference, Transformers
[…]
captionUnknown document class
@@ 1 Introduction
Generative modelling has seen remarkable advances in recent years, with diffusion models achieving state-of-the-art performance across multiple domains, including images, videos, and 3D synthesis (Ho et al., 2020; Dhariwal & Nichol, 2021; Rombach et al., 2022). A key limitation of existing generative frameworks, however, is their reliance on structured output representations, such as pixel grids, which constrain resolution and generalisation across data modalities. In contrast, Implicit Neural Representations (INRs) have emerged as a powerful alternative that parametrises signals as continuous functions (Sitzmann et al., 2020; Mildenhall et al., 2021). By leveraging INRs, generative models can represent complex data distributions at arbitrary resolutions, yet learning expressive distributions over INR parameters remains a fundamental challenge.
@@
Figure 1: Samples from LDMI at multiple resolutions.
A common approach to modelling distributions over INRs is to use hypernetworks (Ha et al., 2017), which generate the weights and biases of an INR conditioned on a latent code (Dupont et al., 2022b; Koyuncu et al., 2023). However, MLP-based hypernetworks suffer from scalability bottlenecks, particularly when generating high-dimensional INRs, as direct parameter regression constrains flexibility and expressiveness. More recently, Transformer-based hypernetworks have been proposed to alleviate these issues, introducing attention mechanisms to efficiently predict INR parameters (Chen & Wang, 2022; Zhmoginov et al., 2022). Nonetheless, existing approaches such as Trans-INR (Chen & Wang, 2022) remain deterministic, limiting their applicability in probabilistic frameworks.
In this work, we introduce a novel framework for INR generation, named Latent Diffusion Models of INRs (LDMI), that incorporates our proposed Hyper-Transformer Decoder (HD), a probabilistic Transformer-based decoder for learning distributions over INR parameters. Our approach combines the strengths of hypernetworks with recent advances in meta-learning for INRs while integrating into latent diffusion-based generative frameworks. Unlike previous works, our HD decoder employs a full Transformer architecture, where a Transformer Encoder processes latent variables, and a Transformer Decoder generates INR parameters via cross-attention. This design enables flexible, probabilistic generation of neural representations across diverse data modalities, surpassing the deterministic limitations of prior Transformer-based hypernetworks (Chen & Wang, 2022; Zhmoginov et al., 2022).
Our proposed framework, LDMI, supports two training paradigms: (i) full training, where the model is trained from scratch alongside a latent diffusion model (LDM; Rombach et al., 2022), and (ii) hyper-transforming, where a pre-trained LDM is adapted by replacing its decoder with the HD, allowing for efficient transfer learning without retraining the entire generative pipeline. This flexibility enables LDMI to scale effectively while leveraging existing pre-trained diffusion models.
@@ 2.1 Latent Diffusion Models
Latent Diffusion Models (LDM; Rombach et al., 2022) are a class of generative models that learn efficient representations by applying diffusion processes in a compressed latent space. Unlike traditional diffusion models that operate directly in high-dimensional data spaces ${\bm{y}}\in\mathbb{R}^{D}$ to approximate $p({\bm{y}})$, LDMs first encode data into a lower-dimensional latent space ${\bm{z}}\in\mathbb{R}^{d}$, where $d\ll D$, using a probabilistic encoder $\mathcal{E}_{\psi}(\cdot)$ to produce rich latent representations $q_{\psi}({\bm{z}}|{\bm{y}})$. These are decoded via $\mathcal{D}_{\lambda}(\cdot)$ to accurately recover data samples by modeling the conditional distribution $p_{\lambda}({\bm{y}}|{\bm{z}})$. A $\beta$-VAE (Higgins et al., 2017) with low $\beta$, a Vector-Quantized VAE (VQ-VAE; van den Oord et al., 2017), or VQGAN (Esser et al., 2021) are suitable choices for learning the structured latent space. Additionally, for image data, perceptual losses can be incorporated to further enhance generation quality (Larsen et al., 2016; Johnson et al., 2016; Hou et al., 2017; Dosovitskiy & Brox, 2016; Hou et al., 2019).
@@
DDPMs (Ho et al., 2020) train using an objective derived from a variational lower bound on the negative log-likelihood, reparameterized for efficiency as a denoising task. This results in a simplified loss that closely resembles denoising score matching (Song et al., 2021b), and is widely used in state-of-the-art models (Saharia et al., 2022; Dhariwal & Nichol, 2021). For latent diffusion, the loss is applied to ${\bm{z}}\sim q_{\psi}({\bm{z}}|{\bm{y}})$, yielding:
$L_{\text{DDPM}}=\mathbb{E}_{{\bm{y}},{\bm{z}},t,\epsilon}\left[\|\epsilon-\epsilon_{\theta}({\bm{z}}_{t},t)\|^{2}\right],$ (1)
where ${\bm{y}}\sim p_{\text{data}}({\bm{y}})$, ${\bm{z}}\sim q_{\psi}({\bm{z}}|{\bm{y}})$, $t\sim\mathcal{U}(1,T)$, and $\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. This objective corresponds to predicting the added noise in a fixed forward process, effectively denoising ${\bm{z}}_{t}$ toward the original latent ${\bm{z}}$.
@@
In DDPM, the reverse posterior density is no longer Markovian and coincides with the inference model proposed later in DDIM (Song et al., 2021a). In DDIM, it is demonstrated that faster sampling can be achieved without retraining, simply by using the posterior approximation for the estimation $\hat{{\bm{z}}}$, which re-defines the generative process as:
@@ 2.2 Probabilistic Implicit Neural Representations
Implicit Neural Representations (INRs; Sitzmann et al., 2020) model continuous functions by mapping coordinates ${\bm{x}}$ to output features ${\bm{y}}$ via a neural network $f_{\Phi}({\bm{y}}|{\bm{x}})$ (Mildenhall et al., 2021; Mescheder et al., 2019; Tancik et al., 2020). Recent generative approaches adapt this idea for flexible conditional generation at arbitrary resolutions (Dupont et al., 2022a; Koyuncu et al., 2023). Unlike models that fit $p({\bm{y}})$ over structured data, INRs approximate the conditional $p({\bm{y}}|{\bm{x}})$, allowing data sampling and uncertainty estimation at any location—especially beneficial for spatially-varying data.
Probabilistic INR generation typically involves meta-learning hidden representations that are mapped to INR parameters $\Phi$. The main challenge lies in learning these representations under uncertainty, with two goals: (i) embedding observed data into a posterior for conditional generation, and (ii) sampling from a prior for synthesis. Koyuncu et al. (2023) address this via a latent variable ${\bm{z}}$ and a variational framework combining a flow-based prior $p_{\theta}({\bm{z}})$ with an encoder $q_{\psi}({\bm{z}}|{\bm{x}},{\bm{y}})$, using an MLP-based hypernetwork to map ${\bm{z}}$ to $\Phi$ (Ha et al., 2017). However, scaling to high-resolution images is limited by the complexity of the INR parameter space. Similarly, Dupont et al. (2022b) use adversarially trained hypernetworks, which improve stability but lack conditional generation and face the same scalability bottlenecks.
Two recent strategies aim to address these issues. Functa (Dupont et al., 2022a; Bauer et al., 2023) learns neural fields per datapoint via optimization of modulation codes (referred to as functas), then trains a generative model on the resulting functaset. Alternatively, Chen et al. (2024) leverage LDMs by pretraining an autoencoder (e.g., $\beta$-VAE or VQ-VAE) with an image renderer incorporated to the decoder, and applying a diffusion model on its latent space.
@@ 2.3 Hyper-Transformers
Hypernetworks (Ha et al., 2017) map latent representations ${\bm{z}}$ to INR parameters $\Phi=g_{\phi}({\bm{z}})$, where $g_{\phi}$ is a learned generator and $f_{\Phi}({\bm{y}}|{\bm{x}})$ defines the INR. While the hypernetwork is shared across the dataset, the INR is instance-specific, generated dynamically by the hypernetwork. This design, however, struggles with large INRs, as generating full parameter vectors introduces capacity and optimization bottlenecks. To overcome this, recent work has proposed Hyper-Transformers (Chen & Wang, 2022; Zhmoginov et al., 2022), which use attention to generate INR parameters in modular chunks. This design avoids flattening and allows scalable, targeted weight generation.
Building on this paradigm, recent work, (HyperDreamBooth; Ruiz et al., 2024) extends this idea by generating LoRA adapters for personalizing diffusion models from few examples. While their focus is efficient adaptation, our framework (LDMI) builds a full generative model over INR parameter space. Our Hyper-Transformer Decoder enables resolution-agnostic generation across modalities (images, 3D fields, climate data), representing a new direction in function-level generative modeling.
Although prior methods succeed in deterministic reconstruction, they do not support probabilistic generation or serve as decoders for latent variable models. Our approach addresses this by using full Transformer architectures conditioned on tensor-shaped latent variables.
@@ 2.4 Function-valued Stochastic Processes
Alternative approaches like Neural Diffusion Processes (NDPs) (Dutordoir et al., 2023) and Simformer (Gloeckler et al., 2024) define distributions over function values, not over network parameters. NDPs denoise function values directly using permutation-invariant attention, mimicking Neural Processes, while Simformer learns a joint diffusion model over data and simulator parameters. In contrast, our method directly generates INR parameters via a hyper-transformer, offering continuous function generation with resolution independence.
@@ 3 Methodology
(a) Encoder.
@@
Figure 2: Overview of the I-VAE. (a) The inference model, trained in the first stage, where an encoder maps data into the variational parameters of the approximate posterior; (b) The full decoder: the HD module transforms the latent variable into the weights and biases of an INR, enabling continuous signal representation.
We introduce Latent Diffusion Models of Implicit Neural Representations (LDMI), a novel family of generative models operating in function space. While previous probabilistic INR approaches have explored latent variable models for INRs (Koyuncu et al., 2023; Dupont et al., 2022b) and two-stage training schemes incorporating diffusion-based priors in the second stage (Park et al., 2023; Dupont et al., 2022a; Bauer et al., 2023), these methods face key limitations. Notably, MLP-based hypernetworks suffer from capacity bottlenecks, while existing frameworks lack a unified and compact approach for handling diverse data modalities.
To address these challenges, we propose the Hyper-Transformer Decoder (HD), which extends the flexibility of hypernetworks with the scalability of Transformer-based architectures. Unlike prior deterministic frameworks (Chen & Wang, 2022), HD introduces a probabilistic formulation, enabling uncertainty modelling and improved generative capacity for INRs.
@@ 3.2 Generative Model
A common approach in prior works (Koyuncu et al., 2023; Dupont et al., 2022b, a) is to model $\Phi$ implicitly using an auxiliary neural network known as a hypernetwork (Ha et al., 2017), typically MLP-based, which modulates the INR parameters based on a lower-dimensional latent representation:
@@
where ${\bm{z}}\in\mathbb{R}^{H_{z}\times W_{z}\times d_{z}}$ is a tensor-shaped latent code with spatial dimensions $H_{z}\times W_{z}$ and channel dimension $d_{z}$. This latent space serves as a compressed representation of the continuous signal. However, MLP-based hypernetworks introduce a fundamental bottleneck: the final layer must output all parameters of the target INR, which leads to scalability issues when modulating high-capacity INRs. As a result, prior works employing MLP-based hypernetworks typically use small INR architectures, often restricted to three-layer MLPs (Koyuncu et al., 2023; Dupont et al., 2022b). Additionally, these works rely on standard MLPs with ReLU or similar activations and often preprocess coordinates using Random Fourier Features (RFF) (Tancik et al., 2020) to capture high-frequency details. However, RFF-based embeddings struggle to generalize to unseen coordinates without careful validation. In contrast, SIREN (Sitzmann et al., 2020) inherently captures the full frequency spectrum by incorporating sinusoidal activations, making it more effective for super-resolution tasks.
However, using SIREN as the INR module introduces further challenges. As commented by their authors, optimizing SIRENs with not carefully chosen uniformly distributed weights yields poor performance both in accuracy and in convergence speed. This issue worsens when the weights are not optimized, but generated by a hypernetwork.
Figure 3: Diagram of the Hyper-Transformer Decoder (HD). The latent variable ${\bm{z}}$ is tokenized and processed by a Transformer Encoder. A Transformer Decoder, initialized with learnable grouped weights $\bar{{\mathbf{W}}}_{l}^{\text{i}}$, cross-attends to the latent tokens to generate the set of grouped weights ${\mathbf{W}}_{l}^{\text{o}}$. The full weight matrices ${\mathbf{W}}_{l}$ are then reconstructed by combining the grouped weights with learnable template weights $\bar{{\mathbf{W}}}_{l}^{\text{b}}$. Biases $\mathbf{b}_{l}$ are learned as global parameters.
To overcome these limitations, we introduce the Hyper-Transformer Decoder (HD), a Transformer-based hypernetwork designed to scale effectively while preserving the flexibility of larger INRs and ensuring proper modulation of SIREN weights.